hallm:

Solve the system of equations using the linear combination method. {4x−3y=12 {7x−3y=3 (its supposed to look like one big { that makes both of the equations go in it Enter your answers in the boxes. x = y =

hallm:



dude:

y is the easiest to get rid of, you need one equation to have a +3 (Multiply one whole equation by -1)

bammbiii5:

\[4x-3y=12 \] and \[7x-3y=3\] Solve x for \[4x-3y=12\] \[4x−3y+3y=12+3y \]\[4x=3y+12 \]\[4x/4 = 3y+12/4\]\[X=3/4y+3\]\[7x−3y=3 \]\[7(3/4y+3)−3y=3\]\[9/4y+21=3\]\[9/4y+21+−21=3+−21\]\[9/4y=−18\]\[9/4y/9/4 = -18/9/4\]\[y=−8 \]\[x=3/4y+3\]\[x=3/4(−8)+3\]\[x=−3 \] \[x=−3 \]\[y=-8\]